5th International Conference on Logic Programming and Nonmonotonic Reasoning LPNMR’99

El Paso, Texas USA, December 2–4, 1999

CALL FOR PAPERS

LPNMR’99 is the fifth in the series of international meetings on logic programming and nonmonotonic reasoning. Four previous meetings were held in Washington, U.S.A., in 1991, in Lisbon, Portugal, in 1993, in Lexington, U.S.A., in 1995, and in Dagstuhl, Germany, in 1997. LPNMR’99 will be coordinated with the ICLP’99, which starts on November 29 in Las Cruces, approximately 40 miles from El Paso.

The proceedings of LPNMR’99 will be published by Springer in the LNCS/LNAI series

(http://www.springer.de/comp/lncs/).

Location

El Paso, Texas is a city of nearly three-quarters of a million people, which sprawls across hundreds of square miles of desert and rambling foothills. The Franklin Mountains, southern tip of the Rockies, slice El Paso nearly in two. With its classic Western geography and tri-cultural history, El Paso makes an interesting place to visit. El Paso is the gateway to Juarez, Mexico (US citizens do not need a visa for a short visit across the border), and to southern New Mexico with such recognized attractions as Fort Bliss, Carlsbad Caverns and White Sands National Memorial. El Paso’s dry, sunny, warm desert climate earned the city the nickname Sun City.

Aim and Scope

The aim of the conference is to facilitate interactions between researchers interested in the design and implementation of logic based programming languages and database systems and researchers who work in the areas of knowledge representation and non-monotonic reasoning.

A non-exhaustive list of topics of interest include:

1. Development and mathematical investigation of logical systems with non-monotonic entailment relations. This includes (but is not limited to)

(a) Extensions of “classical” LPNMR languages by new logical connectives and new inference capabilities such as abduction, reasoning by cases, etc;

(b) Semantics of new and existing languages;

(c) Relationship between various formalisms;

(d) Complexity and expressive power;

(e) Development and implementation of inference mechanisms for LPNMR systems;

(f) Updates and other operations on LPNMR systems;

(g) LPNMR systems with uncertainty.

2. Applications of LPNMR systems.

(a) Methodology of representing knowledge in LPNMR languages. Theory and practice;

(b) LPNMR languages and algorithms in planning, diagnoses, software engineering, and other domains;

(c) Implemented LPNMR systems: Descriptions, Comparisons, Evaluations and Benchmarks.

Submission of Papers

Papers must not exceed twelve (12) pages including references and figures, with a maximum of 38 lines per page and an average of 75 characters per line (corresponding to the LaTeX article style, 12 pt). Further information

is available on the Springer website at http://www.springer.de/comp/lncs/authors.html.

Papers must be written in English and include a cover page containing: title, abstract, keywords, postal and email addresses of all authors, voice and fax number of the contact author. Accepted papers must be presented at the conference. Please consult our web page for detailed submission information.

Important Dates

Friday May 28, 1999: Papers due Monday July 26, 1999: Notification of acceptance/rejection Monday August 23, 1999: Final camera-ready papers due Thursday–Saturday December 2–4, 1999: LPNMR’99

Program Co-Chairs

Michael Gelfond (University of Texas at El Paso, USA) Nicola Leone (Vienna University of Technology, Austria)

Program Committee

Jose Julio Alferes (Universidade de Evora, Portugal) Chitta Baral (University of Texas at El Paso, USA) Nicole Bidoit (Université de Bordeaux 1, France) Jürgen Dix (University of Koblenz, Germany) Thomas Eiter (Vienna University of Technology, Austria) Fangzhen Lin (The Hong Kong University of Science and Technology, China) Jack Minker (University of Maryland, USA) Anil Nerode (Cornell University, USA) Ilkka Niemela (Helsinki University of Technology, Finland) Dino Pedreschi (University of Pisa, Italy) Pasquale Rullo (University of Calabria, Rende, Italy) Chiaki Sakama (Wakayama University, Japan) V.S. Subrahmanian (University of Maryland, USA) Francesca Toni (Imperial College, London, U.K.) Miroslaw Truszczynski (University of Kentucky at Lexington, USA) Hudson Turner (University of Minnesota at Duluth, USA) Moshe Y. Vardi (Rice University, USA) Jia-Huai You (University of Alberta, Canada)

Invited Speakers

Marco Cadoli (Università di Roma ”La Sapienza”, Italy) Vladimir Lifschitz (University of Texas at Austin, USA) David McAllester (AT&T Labs Research, USA) Leora Morgenstern (IBM T.J. Watson Research Center, USA)

Publicity Chair

Gerald Pfeifer (Vienna University of Technology, Austria)

For further details and up-to-date information please check our web page at

http://www.dbai.tuwien.ac.at/lpnmr99/.